# The vast world of quantum advantage

### v1

@@ Acknowledgments:
The authors thank Sergio Boixo, Ryan Babbush, and Stephen Jordan for a careful reading and helpful comments. We would also like to thank Robbie King and Mikhail Lukin for illuminating discussions, and Chi-Yun Cheng for discussions and assistance with the main figure. The authors acknowledge Mauro Schiulaz and other editors from Physical Review X for insightful discussions that led to the inception of this perspective article. H.H. acknowledges support from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator and the Broadcom Innovation Fund. S.C. acknowledges the support from the Center for Ultracold Atoms, an NSF Physics Frontiers Center (Grant No. 2317134). J.P. acknowledges support from the U.S. Department of Energy, Office of Science, Accelerated Research in Quantum Computing, Quantum Utility through Advanced Computational Quantum Algorithms (QUACQ) and Fundamental Algorithmic Research toward Quantum Utility (FAR-Qu), from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator, and from the Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (PHY-2317110).
@@ A.2 Puzzle I: Exponential quantum advantage in analyzing big data?
The construction of efficient sampling access from standard RAM is a well-known technique in randomized algorithms Devroye (1986); Vose (1991). The goal is to sample from a discrete probability distribution where outcome $i$ has probability $p_{k}(i)=|x_{k,i}|^{2}$. The naive approach would require $\mathcal{O}(2^{n})$ time per sample by iterating through all entries, but preprocessing enables much faster sampling.
@@
This solution exemplifies a broader pattern in quantum algorithm analysis. Similar classical techniques have appeared across different fields, including neural quantum states Carleo and Troyer (2017), distributed computing for inner product estimation Buhrman et al. (2010), randomized linear algebra Mahoney (2011), and the dequantization of quantum algorithms Tang (2019). These examples demonstrate how classical probabilistic techniques can sometimes achieve the same performance as quantum algorithms when operating under comparable data-access assumptions.
@@ A.3 Puzzle I: Can quantum systems encode exponential classical information?
The answer to this puzzle is a nuanced “yes,” that reveals a subtle but powerful form of quantum advantage. The apparent conflict between exponential encoding and Holevo’s bound is resolved by understanding that the usefulness of an encoding depends entirely on the task for which the information is used. Holevo’s bound Holevo (1973) states that it is impossible for Bob to learn all $2^{n}$ classical values $x_{i}$ from the following $n$-qubit state:
@@ C.1 Overview and Context
The low-weight Pauli propagation algorithm Aharonov et al. (2023); Shao et al. (2023); Nemkov et al. (2023); Begušić et al. (2024, 2023); Fontana et al. (2023); Rudolph et al. (2023); Angrisani et al. (2024) provides a powerful classical heuristic for simulating quantum circuits. The key idea is to truncate high-weight Pauli terms during the simulation, maintaining only the most important contributions. Consider a quantum circuit with $L$ layers:
@@ C.3 Theoretical Guarantees
When each circuit layer $U_{j}$ is drawn from a locally scrambling distribution Kuo et al. (2020); Hu et al. (2023); Caro et al. (2023); Gibbs et al. (2024); Huang et al. (2022b), which corresponds to a probability distribution that remains invariant under single-qubit Clifford rotations, Angrisani et al. (2024) has established the following theorem to provide a rigorous guarantee for the low-weight Pauli propagation algorithm. Intuitively, the locally scrambling property means each layer is generic on a local scale and mixes the local basis.
@@ C.4 Applicability and Scope
    Brickwork circuits of $SU(4)$ gates in 1D, 2D, and 3D Zhang et al. (2024); Napp (2022); Braccia et al. (2024).
@@
    Circuits with universal single-qubit rotations followed by entangling Clifford gates Letcher et al. (2023).
@@
    Quantum Convolutional Neural Networks without feed-forward Pesah et al. (2021).
@@ Appendix D Classical hardness in predicting quantum advantages
Step 1: Random Circuit Construction. Let $U=U_{L}\cdots U_{1}$ be an $n$-qubit random quantum circuit where each $U_{j}$ is a layer of random two-qubit gates, and the total depth $L$ can be chosen arbitrarily as long as $L$ is polynomial in $n$ and $m$. By choosing $L$ to be at least linear in $n$ with a sufficiently large constant, it is known that $U$ forms an $\epsilon$-approximate unitary $2$-design with $\epsilon=\mathcal{O}(1/2^{n})$ Harrow and Low (2009); Brandao et al. (2016); Schuster et al. (2024). This means the distribution over the random unitary $U$ nearly matches the Haar measure over all $n$-qubit unitaries $\mathrm{SU}(2^{n})$ up to the second moment.
@@
    from the standard Weingarten calculation for Haar-random unitaries Harrow and Low (2009); Brandao et al. (2016). Here, $\delta_{P_{i},P^{\prime\prime}_{i^{\prime\prime}}}$ is $1$ if the two $n$-qubit Pauli operators $P_{i},P^{\prime\prime}_{i^{\prime\prime}}$ are the same and 0 otherwise. If $i^{\prime}$ is not acted on by the two-qubit gate $V$, then because $P^{\prime}_{i^{\prime}}$ is traceless, we have

### v2

@@ Acknowledgments:
The authors thank Sergio Boixo, Ryan Babbush, Edward Farhi, Sam Gutmann, and Stephen Jordan for a careful reading and helpful comments. We would also like to thank Robbie King and Mikhail Lukin for illuminating discussions, and Chi-Yun Claudia Cheng for discussions and assistance regarding the main figure. The authors acknowledge Mauro Schiulaz and other editors from Physical Review X for insightful discussions that led to the inception of this perspective article. H.H. acknowledges support from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator and the Broadcom Innovation Fund. S.C. acknowledges the support from the Center for Ultracold Atoms, an NSF Physics Frontiers Center (Grant No. 2317134). J.P. acknowledges support from the U.S. Department of Energy, Office of Science, Accelerated Research in Quantum Computing, Quantum Utility through Advanced Computational Quantum Algorithms (QUACQ) and Fundamental Algorithmic Research toward Quantum Utility (FAR-Qu), from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator, and from the Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (PHY-2317110).
@@ A.2 Puzzle I: Exponential quantum advantage in analyzing big data?
The construction of efficient sampling access from standard RAM is a well-known technique in randomized algorithms [118, 119]. The goal is to sample from a discrete probability distribution where outcome $i$ has probability $p_{k}(i)=|x_{k,i}|^{2}$. The naive approach would require $\mathcal{O}(2^{n})$ time per sample by iterating through all entries, but preprocessing enables much faster sampling.
@@
This solution exemplifies a broader pattern in quantum algorithm analysis. Similar classical techniques have appeared across different fields, including neural quantum states [76], distributed computing for inner product estimation [120], randomized linear algebra [121], and the dequantization of quantum algorithms [5]. These examples demonstrate how classical probabilistic techniques can sometimes achieve the same performance as quantum algorithms when operating under comparable data-access assumptions.
@@ A.3 Puzzle I: Can quantum systems encode exponential classical information?
The answer to this puzzle is a nuanced “yes,” that reveals a subtle but powerful form of quantum advantage. The apparent conflict between exponential encoding and Holevo’s bound is resolved by understanding that the usefulness of an encoding depends entirely on the task for which the information is used. Holevo’s bound [7] states that it is impossible for Bob to learn all $2^{n}$ classical values $x_{i}$ from the following $n$-qubit state:
@@ C.1 Overview and Context
The low-weight Pauli propagation algorithm [33, 34, 35, 36, 37, 38, 39, 40, 41] provides a powerful classical heuristic for simulating quantum circuits. The key idea is to truncate high-weight Pauli terms during the simulation, maintaining only the most important contributions. Consider a quantum circuit with $L$ layers:
@@ C.3 Theoretical Guarantees
When each circuit layer $U_{j}$ is drawn from a locally scrambling distribution [122, 123, 124, 125, 126], which corresponds to a probability distribution that remains invariant under single-qubit Clifford rotations, [40] has established the following theorem to provide a rigorous guarantee for the low-weight Pauli propagation algorithm. Intuitively, the locally scrambling property means each layer is generic on a local scale and mixes the local basis.
@@ C.4 Applicability and Scope
    Brickwork circuits of $SU(4)$ gates in 1D, 2D, and 3D [127, 128, 129].
@@
    Circuits with universal single-qubit rotations followed by entangling Clifford gates [130].
@@
    Quantum Convolutional Neural Networks without feed-forward [131].
@@ Appendix D Classical hardness in predicting quantum advantages
Step 1: Random Circuit Construction. Let $U=U_{L}\cdots U_{1}$ be an $n$-qubit random quantum circuit where each $U_{j}$ is a layer of random two-qubit gates, and the total depth $L$ can be chosen arbitrarily as long as $L$ is polynomial in $n$ and $m$. By choosing $L$ to be at least linear in $n$ with a sufficiently large constant, it is known that $U$ forms an $\epsilon$-approximate unitary $2$-design with $\epsilon=\mathcal{O}(1/2^{n})$ [132, 133, 19]. This means the distribution over the random unitary $U$ nearly matches the Haar measure over all $n$-qubit unitaries $\mathrm{SU}(2^{n})$ up to the second moment.
@@
    from the standard Weingarten calculation for Haar-random unitaries [132, 133]. Here, $\delta_{P_{i},P^{\prime\prime}_{i^{\prime\prime}}}$ is $1$ if the two $n$-qubit Pauli operators $P_{i},P^{\prime\prime}_{i^{\prime\prime}}$ are the same and 0 otherwise. If $i^{\prime}$ is not acted on by the two-qubit gate $V$, then because $P^{\prime}_{i^{\prime}}$ is traceless, we have